OpenStudy (nerdgamer_):

2 and one-third plus 3 and one-half is equal to what number? 5 and two-fifths 5 and two-thirds 5 and one-half 5 and five-sixths

OpenStudy (nerdgamer_):

`Please Help I Will Give A Medal & A Fan To Whoever Helps Me!! :)...........................................I really need help!`

OpenStudy (nerdgamer_):

@allopersonwhat @Skyz @WHAT?! @genius12 @10_4 @zubhanwc3 @phi @thomaster @ElloLovee @Ambbiiee @Luigi0210 @dape @potato_donuts @CGGURUMANJUNATH @forevershorty @shaik0124 @texaschic101 @annie_r_u_otay @primeralph @Emmalineh123 @ehuman Halp!!

OpenStudy (nerdgamer_):

`Please help me!!!`

OpenStudy (forevershorty):

it's A

OpenStudy (zubhanwc3):

o misread o.O

OpenStudy (nerdgamer_):

`Do you know why its A?? @forevershorty `

OpenStudy (zubhanwc3):

its d

OpenStudy (nerdgamer_):

`Do you know why?`

OpenStudy (shaik0124):

five and five -sixths

OpenStudy (forevershorty):

I suck at math listen to them xP

OpenStudy (zubhanwc3):

\[2\frac{ 1 }{ 3 }+ 3\frac{ 1 }{ 2 }\]

OpenStudy (nerdgamer_):

`I cant trust everybody!`

OpenStudy (nerdgamer_):

ohh ok lol

OpenStudy (anonymous):

its 35/6 which could be converted into 5 and 5/6.

OpenStudy (shaik0124):



OpenStudy (zubhanwc3):

it becomes \[5(\frac{ 1 }{ 2 }+\frac{ 1 }{ 3 })\] add the two and u got \[5\frac{ 5 }{ 6 }\]

OpenStudy (nerdgamer_):

\(Everybody is saying somthin different!!!!\)

OpenStudy (zubhanwc3):

no we are not, we are using different methods.

OpenStudy (nerdgamer_):

okay I belive you @zubhanwc3 &@ElloLovee

OpenStudy (anonymous):

The answer is D ! Alright ? Listen to me!

OpenStudy (anonymous):

im really good with fractions.

OpenStudy (nerdgamer_):

`Okay thxs!!!` @ElloLovee & @zubhanwc3